OpenStudy (anonymous):

Check all polynomials that are perfect square trinomials: 9x2 − 6x + 1 36b2 − 24b + 8 16x2 + 24x + 9 4a2 − 10a + 25

OpenStudy (anonymous):

The first term has to be a perfect square, the last term has to be a perfect square, and the middle term has to be twice the product of the first and last terms square root

OpenStudy (anonymous):

1 and 8 aren’t perfect squares. The square root of 9=?

OpenStudy (anonymous):

so is it 9x2-6x + 1 ?

OpenStudy (anonymous):

No, the square root of 9x2= 18 not 24, so which one is it?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

Ok the first 3 are wrong because they do not follow the rules that i posted above. The last one is the only perfect square trinomial because \[\sqrt{4}=2, \sqrt{25}=5, 5\times2=10\]

OpenStudy (anonymous):

A. \[\sqrt{9}=3, \sqrt{1}=1, 1\times2=2\] B. \[\sqrt{36}=6, \] and 8 cant be squared C. \[\sqrt{16}=4, \sqrt{9}=3, 3\times2=6\]

OpenStudy (anonymous):

so do i do check the last one ?

OpenStudy (anonymous):

Yeah \[4a^2-10a+25\]

OpenStudy (anonymous):

okayy thankyouu